# MGMCpaper.bib

### v2

@@ 2.1 Sampling Hilbert space-valued Gaussian random variables
Let $(\Omega,\mathcal{\mathscr{F}},\mathbb{P})$ be a probability space and $(H,\langle\cdot,\cdot\rangle_{H})$ a separable Hilbert space. Let $\mathcal{A}^{-1}\colon H\to H$ be a self-adjoint positive definite operator. Then we aim to generate samples following the Gaussian distribution $\mathcal{N}(0,\mathcal{A}^{-1})$ with covariance operator $\mathcal{A}^{-1}$, where without loss of generality we assume that the mean is zero. An important example of $\mathcal{A}^{-1}$ is the Matérn class of covariance operators, which is given by suitable inverse powers of shifted Laplace operators, such that $\mathcal{A}^{-1}=(-\Delta+\kappa^{2}I)^{-\alpha}$ for $\alpha>0$ [whittle1954stationary, whittle1963stochastic, LindgrenEtAl.F_2011_ExplicitLinkGaussian, BolinEtAl.D_2020_NumericalSolutionFractional].
Note that for $\mathcal{N}(0,\mathcal{A}^{-1})$ to be a distribution of an $H$-valued random variable, $\mathcal{A}^{-1}$ must necessarily be a trace class operator (see e.g., [DaPrato.D_2014_book, Proposition 2.16]). This implies the compactness of $\mathcal{A}^{-1}$, which naturally gives us connections to abstract partial differential equations (PDEs) given by unbounded operators with compact inverse, and their discretisations. More precisely, we consider the weak formulation of abstract PDEs on the Cameron–Martin space and its discrete counterparts, which gives rise to the linear operator $\mathcal{A}$ and corresponding discretisation matrices. If $\mathcal{A}^{-1}$ is compact but not trace-class, it is still possible to define distributions of real-valued random variables that describe observations of an underlying Gaussian random field in a bigger space than $H$; this is explained in more detail in Remark 2.3 below.
2.2 Linear Bayesian inverse problem
@@
where $\mathcal{B}^{*}\colon\mathbb{R}^{\beta}\to H$ is the (Hilbert space) adjoint operator of $\mathcal{B}$; see for example [Hairer.M_etal_2005_AnalysisSPDEsArising, Lemma 4.3] together with [Sullivan.TJ_2015_IntroductionUncertaintyQuantification, Chapter 2]. Further, define $f\in H$ as
@@
We are interested in generating samples from the posterior distribution $\mathcal{N}(\mu,\widetilde{\mathcal{A}}^{-1})$ and this paper focuses on the multigrid Monte Carlo sampler, which was introduced in [goodman1989multigrid] and is described in Section 3 below. The generated samples can be used to compute the expected value of a quantity of interest which is a functional $\mathcal{F}:H\rightarrow\mathbb{R}$. In the following we assume that $\mathcal{F}$ is a bounded linear functional of the form
@@ 2.3 Discretisation
We consider the vector subspace $V:=D(\mathcal{A}^{1/2})\subset H$ with $D(\mathcal{A}^{1/2})$ denoting the domain of the operator $\mathcal{A}^{1/2}$ in $H$. Here, the square-root $\mathcal{A}^{1/2}$ of $\mathcal{A}$ is defined spectrally. Because $\mathcal{A}^{1/2}$ is positive definite, the bilinear form $a(\cdot,\cdot)\colon V\times V\to\mathbb{R}$ defined via
@@
is an inner product on $V$ and thus, $(V,a(\cdot,\cdot))$ forms a Hilbert space; see e.g., [Sell.G_You_2013_book, Section 3.7]. As usual, we let $\|v\|_{V}:=a(v,v)$ and note that then $\|v\|_{H}\leq c\|v\|_{V}$ for $v\in V$ with some constant $c>0$. The space $D(\mathcal{A}^{1/2})$ is identical to the image of $\mathcal{A}^{-1/2}$ on $H$, which is often referred to as the Cameron–Martin space [DaPrato.G_2006_IntroductionInfiniteDimensionalAnalysis, Lifshits.M.A_2012_Lectures_on_Gaussian].
To discretise (2.1)–(2.3) we introduce a hierarchy $V_{0}\subset\dotsb\subset V_{\ell}\subset\dotsb\subset V$ of nested finite dimensional subspaces with dimensions $n_{\ell}$ for $\ell\geq 0$. These subspaces are typically finite element spaces and for each $\ell$ we choose a basis $\{\phi_{j}^{\ell}\}_{j=1,\dots,n_{\ell}}$ of $V_{\ell}$. In each subspace $V_{\ell}$ we observe the quantity
@@
For example, when $H=L^{2}(D)$ for a suitable domain $D\subset\mathbb{R}^{d}$ and $V_{\ell}$ is a finite element space with shape-regular mesh, then Assumption 2.1 is satisfied with $\Phi(\ell)=h_{\ell}^{d/2}$; see [Hackbusch.W_2017_book_elliptic_2nd, Theorem 8.76].
@@
where $\{e_{j}^{\ell}\}_{j=1,2,\dots,n_{\ell}}$ is the canonical basis of $\mathbb{R}^{n_{\ell}}$. The (dual) vector-representation $f_{\ell}\in\mathbb{R}^{n_{\ell}}$ of $f$ in (2.4) is given by
$(f_{\ell})_{j}:=\langle f,\phi_{j}^{L}\rangle_{H}=y^{\top}\Gamma^{-1}\mathcal{B}\phi_{j}^{L}.$ (2.10)
@@
while the prior on $\mathbb{R}^{n_{\ell}}$ is chosen to be $v_{\ell}\sim\mathcal{N}(0,A_{\ell}^{-1})$. Finally, denote by $B_{\ell}$ the matrix representation of the mapping $\mathcal{B}P_{\ell}\colon\mathbb{R}^{n_{\ell}}\to\mathbb{R}^{\beta}$ with respect to the canonical bases of $\mathbb{R}^{n_{\ell}}$ and $\mathbb{R}^{\beta}$, i.e.
@@
which are the discrete versions of (2.2) and (2.3) respectively. In (2.13) we used the Sherman–Morrison–Woodbury formula to express $\widetilde{A}_{\ell}^{-1}$ as a low-rank update of $A_{\ell}^{-1}$; see for example [Stuart.A_2010_Acta_Inverse, Theorem 6.20, (2.16)] together with [Horn.R_Johnson_book_2013_2nd, Chapter 0].
@@
    for a decreasing function $\Psi\colon\mathbb{N}_{0}\to[0,\infty)$.
@@
Let $\mathcal{A}:=-\Delta+\kappa^{2}I$, where $\Delta$ is the Laplace operator with zero-Dirichlet boundary condition and $\kappa^{2}>0$ is a constant. Let $D\subset\mathbb{R}^{d}$ be a bounded and convex domain, and suppose that $(V_{\ell})_{\ell\geq 0}$ is a family of continuous piecewise linear finite element spaces defined on a regular family of triangulations with mesh sizes $h_{\ell}$. Then, it is well known that Assumption 2.2 is satisfied with $W=H^{2}(D)$ and $\Psi(\ell)\asymp h_{\ell}^{2}$; see, e.g., [Hackbusch.W_2017_book_elliptic_2nd, Theorem 9.24] and [Quarteroni.A_Vali_book_1994_NAPDE, Theorem 3.4.2], respectively.
Under Assumption 2.2, the $H$-norm error $\|u-u_{\ell}\|_{H}$ of approximating $u$ with $u_{\ell}$ decays at the rate $\Psi(\ell)$ as stated in the following proposition (for a proof see, e.g., [Hackbusch.W_2017_book_elliptic_2nd, Proof of Theorem 8.65]).
@@
Let $u\in V$ satisfy (2.17) and let Assumption 2.2 hold. Suppose furthermore that $u_{\ell}\in V_{\ell}$ satisfies
@@
To realise such random variables, we follow [DaPrato.D_2014_book, Section 4.1.2] and let $(\varphi_{j})_{j\geq 1}$ be a complete $H$-orthonormal system (ONS) consisting of eigenfunctions of $\mathcal{A}^{-1}$ with corresponding eigenvalues $(1/\lambda_{j})_{j\geq 1}$. Then, $(\varphi_{j}/\sqrt{\lambda_{j}})_{j\geq 1}$ is a complete ONS of $V=D(\mathcal{A}^{1/2})\subset H$. For each $\chi\in H$, we define $(v,\chi)$ as the $L^{2}(\Omega)$-limit of the series
@@
the series in (2.20) converges the resulting random variable $(v,\chi)$ has the desired mean and covariance in (2.18)–(2.19). Moreover, following again [DaPrato.D_2014_book] the random field $v$ can be directly constructed on a suitably large Hilbert space without taking the coupling with $\chi\in H$.
@@
$\mathbb{E}[(\widetilde{v},\chi)]=$ $\ \langle\chi,\mu\rangle_{H}\quad$ for $\text{for $\chi\in H$},$ (2.21)
@@
For the remainder of this paper, in particular to show $L$-independent convergence, we will only rely on the fact that $\mathcal{A}^{-1}$ is compact not that it is trace-class.
@@ 3.1 Multigrid Monte Carlo
Given some sufficiently large value $L\in\mathbb{N}$, our goal is to sample from the multivariate normal distribution $\mathcal{N}(\mu_{L},\widetilde{A}_{L}^{-1})$ with mean $\mu_{L}:=\widetilde{A}_{L}^{-1}f_{L}$ defined in (2.12) and covariance $\widetilde{A}_{L}^{-1}$ given in (2.13). Recall that $f_{L}\in\mathbb{R}^{n_{L}}$ and $\widetilde{A}_{L}$ is a symmetric $n_{L}\times n_{L}$ matrix. Starting from some initial state $\theta_{L}^{(0)}\in\mathbb{R}^{n_{L}}$, the Multigrid Monte Carlo update in Alg. 1 below will generate a Markov chain $\theta_{L}^{(0)},\theta_{L}^{(1)},\theta_{L}^{(2)},\dots$ with $\theta_{L}^{(k+1)}=\mathrm{MGMC}_{L}(\widetilde{A}_{L},f_{L},\theta_{L}^{(k)})$ which converges to the target distribution $\mathcal{N}(\mu_{L},\widetilde{A}_{L}^{-1})$ in a sense that will be made precise in Section 4.3.
Multigrid Monte Carlo introduces a hierarchy of coarser levels $L-1,L-2,\dots,0$, associated with vector spaces of dimension $n_{L}>n_{L-1}>\dotsb>n_{1}>n_{0}$. This corresponds to the hierarchy $V_{0}\subset V_{1}\subset\dots\subset V_{L-1}\subset V_{L}$ of nested function spaces introduced in Section 2.3 in the sense that the vector space $\mathbb{R}^{n_{\ell}}$ contains the degrees-of-freedom vectors of the function space $V_{\ell}$. In [goodman1989multigrid], the method was introduced for more general distributions, but we restrict ourselves to the Gaussian setting here. On the other hand, our Alg. 1 generalises [goodman1989multigrid] in that it allows arbitrary linear operators for transfering samples between levels of the hierarchy, not just piecewise constant prolongation. As in the standard multigrid algorithm for solving linear systems, we introduce a prolongation matrix $I_{\ell-1}^{\ell}$ defined in (2.14); this matrix has full column rank for all levels $\ell$. The corresponding restriction matrix $I_{\ell}^{\ell-1}=(I_{\ell-1}^{\ell})^{\top}$ is the transpose of the prolongation matrix. Coarse level matrices are then recursively constructed via the so-called Galerkin triple-product
@@ 3.1.1 Heuristic explanation
9:         Define $f_{\ell-1}:=I_{\ell}^{\ell-1}(f_{\ell}-\widetilde{A}_{\ell}\theta_{\ell,\nu_{1}})$triangleright Restriction
@@
12:              $\psi_{\ell-1}^{(m+1)}:=\mathrm{MGMC}_{\ell-1}(\widetilde{A}_{\ell-1},f_{\ell-1},\psi_{\ell-1}^{(m)})$triangleright Recursive call to $\mathrm{MGMC}_{\ell-1}$
@@
14:         $\theta_{\ell,\nu_{1}+1}:=\theta_{\ell,\nu_{1}}+I_{\ell-1}^{\ell}\psi_{\ell-1}^{(\gamma_{\ell})}$triangleright Prolongation
@@
16:              $\theta_{\ell,\nu_{1}+2+j}:=\mathrm{RandomSmoother}(\widetilde{A}_{\ell},\widetilde{M}_{\ell}^{\mathrm{post}},f_{\ell},\theta_{\ell,\nu_{1}+1+j})$triangleright Random post-smoothing (Alg. 2)
@@
    If the initial state $\theta_{L}^{0}\sim\mathcal{N}(\mu_{L}^{0},(\widetilde{A}_{L}^{0})^{-1})$ is drawn from a multivariate normal distribution with mean $\mu_{L}^{0}=(\widetilde{A}_{L}^{0})^{-1}f_{L}^{0}$ and covariance $(\widetilde{A}_{L}^{0})^{-1}$, the Markov chain $\theta_{L}^{(0)},\theta_{L}^{(1)},\theta_{L}^{(2)},\dots$ generated by $\textrm{MGMC}(\widetilde{A}_{L},f_{L},\cdot)$ converges to the target distribution $\mathcal{N}(\mu_{L},\widetilde{A}_{L}^{-1})$ in the sense that $\theta_{L}^{(k)}\sim\mathcal{N}(\mu_{L}^{k},(\widetilde{A}_{L}^{k})^{-1})$ with $\mu_{L}^{k}\rightarrow\mu_{L}$ and $\widetilde{A}_{L}^{k}\rightarrow\widetilde{A}_{L}$. The rate of convergence is grid-independent for both the mean and the covariance (see Section 4.3).
@@
The intuitive explanation for the success of MGMC is the same as for multigrid solvers: by using a hierarchy of levels, the samples are updated on all length scales simultaneously and this incurs only a small, grid independent overhead. This is in stark contrast to a Gibbs sampler, which only updates the samples locally.

### v3

@@ 2.1 Sampling Hilbert space-valued Gaussian random variables
Let $(\Omega,\mathcal{\mathscr{F}},\mathbb{P})$ be a probability space and $(H,\langle\cdot,\cdot\rangle_{H})$ a separable Hilbert space over the real numbers $\mathbb{R}$. Let $\mathcal{A}^{-1}\colon H\to H$ be a self-adjoint positive definite operator. Then we aim to generate samples following the Gaussian distribution $\mathcal{N}(0,\mathcal{A}^{-1})$ with covariance operator $\mathcal{A}^{-1}$, where without loss of generality we assume that the mean is zero. An important example of $\mathcal{A}^{-1}$ is the Matérn class of covariance operators, which is given by suitable inverse powers of shifted Laplace operators, such that $\mathcal{A}^{-1}=(-\Delta+\kappa^{2}I)^{-\alpha}$ for $\alpha>0$ [undefaab, undefaac, undefb, undefaad].
Note that for $\mathcal{N}(0,\mathcal{A}^{-1})$ to be a distribution of an $H$-valued random variable, $\mathcal{A}^{-1}$ must necessarily be a trace class operator (see e.g., [undefaae, Proposition 2.16]). This implies compactness of $\mathcal{A}^{-1}$, which provides natural connections to abstract partial differential equations (PDEs), i.e., unbounded operators with compact inverse, and their discretisations. More precisely, we consider the weak formulation of abstract PDEs on the Cameron–Martin space and its discrete counterparts, which gives rise to the linear operator $\mathcal{A}$ and corresponding discretisation matrices. If $\mathcal{A}^{-1}$ is compact but not trace-class, it is still possible to define distributions of real-valued random variables that describe observations of an underlying Gaussian random field in a bigger space than $H$; this is explained in more detail in Remark 2.3 below and most of our numerical experiments are in this setting.
2.2 Linear Gaussian Bayesian inverse problem
@@
where $\mathcal{B}^{*}\colon\mathbb{R}^{\beta}\to H$ is the (Hilbert space) adjoint operator of $\mathcal{B}$; see for example [undefaaf, Lemma 4.3] together with [undefaag, Chapter 2]. In the finite-dimensional case, it is straightforward linear algebra. Further, define $f\in H$ as
@@
We are interested in generating samples from the posterior distribution $\mathcal{N}(\mu,\widetilde{\mathcal{A}}^{-1})$ and this paper focuses on the multigrid Monte Carlo sampler, which was introduced in [undef] and is described in Section 3 below. The generated samples can be used to sample from the probability distribution of a quantity of interest, which is a functional $\mathcal{F}:H\rightarrow\mathbb{R}$. In the following we assume that $\mathcal{F}$ is a bounded linear functional of the form
@@
While for linear $\mathcal{F}$, considered here, the expectation and covariance of the Gaussian observable can be computed using linear algebra, this is no longer the case for non-linear functionals.
@@ 2.3 Discretisation
We consider the vector subspace $V:=D(\mathcal{A}^{1/2})\subset H$ with $D(\mathcal{A}^{1/2})$ denoting the domain of the operator $\mathcal{A}^{1/2}$ in $H$. Here, the square-root $\mathcal{A}^{1/2}$ of $\mathcal{A}$ is defined spectrally; see [undefaah, Section 3.7] for details. Because $\mathcal{A}^{1/2}$ is positive definite, the bilinear form $a(\cdot,\cdot)\colon V\times V\to\mathbb{R}$ defined via
@@
is an inner product on $V$ and thus, $(V,a(\cdot,\cdot))$ forms a Hilbert space over $\mathbb{R}$; see e.g., [undefaah, Section 3.7]. As usual, we let $\|v\|_{V}:=a(v,v)$ and note that then $\|v\|_{H}\leq c\|v\|_{V}$ for $v\in V$ with some constant $c>0$. The space $D(\mathcal{A}^{1/2})$ is identical to the image of $\mathcal{A}^{-1/2}$ on $H$, which is often referred to as the Cameron–Martin space [undefaai, undefaaj].
To discretise (2.1)–(2.3) we introduce a hierarchy $V_{0}\subset\dotsb\subset V_{\ell}\subset\dotsb\subset V$ of nested finite dimensional subspaces with dimensions $n_{\ell}$ for $\ell\geq 0$. These subspaces are typically finite element (FE) spaces on grids with mesh size $h_{\ell}$ and for each $\ell$ we choose a basis $\{\phi_{j}^{\ell}\}_{j=1,\dots,n_{\ell}}$ of $V_{\ell}$. In each subspace $V_{\ell}$ we observe the quantity
@@
For example, when $H=L^{2}(D)$ for a suitable domain $D\subset\mathbb{R}^{d}$ and $V_{\ell}$ is a FE space with shape-regular mesh, then Assumption 1 is satisfied with $\Phi(\ell)=h_{\ell}^{d/2}$; see [undefaak, Theorem 8.76]. In the language of FE methods, Assumption 1 requires that the mass matrix for the basis functions $\{\phi_{j}^{\ell}\}$ is uniformly well-conditioned, which is a common assumption. For more details on FE methods, such as the construction of FE spaces, we refer the reader to [undefaal].
@@
where $\{e_{j}^{\ell}\}_{j=1,2,\dots,n_{\ell}}$ is the canonical basis of $\mathbb{R}^{n_{\ell}}$. The matrix $A_{\ell}$ is commonly referred to as the stiffness matrix in the FE literature. The (dual) vector-representation $f_{\ell}\in\mathbb{R}^{n_{\ell}}$ of $f$ in (2.4) is given by
$(f_{\ell})_{j}:=\langle f,\phi_{j}^{\ell}\rangle_{H}=y^{\top}\Gamma^{-1}\mathcal{B}\phi_{j}^{\ell}.$ (2.10)
@@
while the prior on $\mathbb{R}^{n_{\ell}}$ is chosen to be $v_{\ell}\sim\mathcal{N}(0,A_{\ell}^{-1})$. Finally, denote by $B_{\ell}$ the transpose of the matrix representation of the mapping $\mathcal{B}P_{\ell}\colon\mathbb{R}^{n_{\ell}}\to\mathbb{R}^{\beta}$ with respect to the canonical bases of $\mathbb{R}^{n_{\ell}}$ and $\mathbb{R}^{\beta}$, i.e.
@@
which are the discrete versions of (2.2) and (2.3) respectively. In (2.13) we used the Sherman–Morrison–Woodbury formula to express $\widetilde{A}_{\ell}^{-1}$ as a low-rank update of $A_{\ell}^{-1}$; see for example [undefaam, Theorem 6.20, (2.16)] together with [undefaan, Chapter 0].
@@
    for a decreasing function $\Psi\colon\mathbb{N}_{0}\to[0,\infty)$ with $\Psi(\ell)\rightarrow 0$ as $\ell\rightarrow\infty$.
@@
Let $\mathcal{A}:=-\Delta+\kappa^{2}I$, where $\Delta$ is the Laplace operator with zero-Dirichlet boundary condition and $\kappa^{2}>0$ is a constant. Let $D\subset\mathbb{R}^{d}$ be a bounded and convex domain, and suppose that $(V_{\ell})_{\ell\geq 0}$ is a family of continuous, piecewise linear FE spaces defined on a regular family of triangulations with mesh sizes $h_{\ell}$. Then, it is well known that Assumption 2 is satisfied with $W=H^{2}(D)$ and $\Psi(\ell)\asymp h_{\ell}^{2}$; see, e.g., [undefaak, Theorem 9.24] and [undefaao, Theorem 3.4.2], respectively.
While Assumption 2 is natural for Matérn-like distributions on finite domains as in Example 2.1, it is more general in that it also covers non-stationary processes similar to the ones described in [undefaap]. To see this consider, for example, a precision operator of the form $\mathcal{A}=-\nabla\cdot(K(x)\nabla\cdot)+\kappa(x)^{2}I$ where $K(x)$ and $\kappa(x)$ are sufficiently smooth, spatially varying fields, respectively matrix- and real-valued with $K(x)$ symmetric positive definite at every point $x\in D$.
Under Assumption 2, the $H$-norm error $\|u-u_{\ell}\|_{H}$ of approximating $u$ with $u_{\ell}$ decays at the rate $\Psi(\ell)$ as stated in the following proposition (for a proof see, e.g., [undefaak, Proof of Theorem 8.65]).
@@
Let $u\in V$ satisfy (2.17) and let Assumption 2 hold. Suppose furthermore that $u_{\ell}\in V_{\ell}$ satisfies
@@
To realise such random variables, we follow [undefaae, Section 4.1.2] and let $(\varphi_{j})_{j\geq 1}$ be a complete $H$-orthonormal system (ONS) consisting of eigenfunctions of $\mathcal{A}^{-1}$ with corresponding eigenvalues $(1/\lambda_{j})_{j\geq 1}$. Then, $(\varphi_{j}/\sqrt{\lambda_{j}})_{j\geq 1}$ is a complete ONS of $V=D(\mathcal{A}^{1/2})\subset H$. For each $\chi\in H$, we define $(v,\chi)$ as the $L^{2}(\Omega)$-limit of the series
@@
the series in (2.20) converges the resulting random variable $(v,\chi)$ has the desired mean and covariance in (2.18)–(2.19). Moreover, following again [undefaae] the random field $v$ can be directly constructed on a suitably large Hilbert space without taking the coupling with $\chi\in H$.
@@
$\mathbb{E}[(\widetilde{v},\chi)]=$ $\ \langle\chi,\mu\rangle_{H}\quad$ $\text{for $\chi\in H$},$ (2.21)
@@
For the remainder of this paper, in particular to show that convergence rates are independent of the mesh-size $h_{\ell}$, we will only rely on the fact that $\mathcal{A}^{-1}$ is compact not that it is trace-class.
@@ 3.1 Multigrid Monte Carlo
Given some sufficiently large value $L\in\mathbb{N}$, our goal is to sample from the multivariate normal distribution $\mathcal{N}(\mu_{L},\widetilde{A}_{L}^{-1})$ with mean $\mu_{L}:=\widetilde{A}_{L}^{-1}f_{L}$ defined in (2.12) and covariance $\widetilde{A}_{L}^{-1}$ given in (2.13). Recall that $f_{L}\in\mathbb{R}^{n_{L}}$ and $\widetilde{A}_{L}$ is a symmetric $n_{L}\times n_{L}$ matrix. Starting from some initial state $\theta_{L}^{(0)}\in\mathbb{R}^{n_{L}}$, the Multigrid Monte Carlo update in Alg. 1 below will generate a Markov chain $\theta_{L}^{(0)},\theta_{L}^{(1)},\theta_{L}^{(2)},\dots$ with $\theta_{L}^{(k+1)}=\mathrm{MGMC}_{L}(\widetilde{A}_{L},f_{L},\theta_{L}^{(k)},\xi^{(k)})$ which converges to the target distribution $\mathcal{N}(\mu_{L},\widetilde{A}_{L}^{-1})$ in a sense that will be made precise in Section 4.3. Here, $\xi^{(k)}$ is a collection of random variables independent of everything before iteration $k$. To simplify notation, we will omit the dependence of $\mathrm{MGMC}_{L}$ on the random variable $\xi^{(k)}$ in the following.
Multigrid Monte Carlo introduces a hierarchy of coarser levels $L-1,L-2,\dots,0$, associated with vector spaces of dimension $n_{L}>n_{L-1}>\dotsb>n_{1}>n_{0}$. This corresponds to the hierarchy $V_{0}\subset V_{1}\subset\dots\subset V_{L-1}\subset V_{L}$ of nested function spaces introduced in Section 2.3 in the sense that the vector space $\mathbb{R}^{n_{\ell}}$ contains the degrees-of-freedom vectors of the function space $V_{\ell}$. In [undef], the method was introduced for more general distributions. While the authors conclude that MGMC does not overcome critical slowing down for theories that undergo a second order phase transition, such as the Ising model in higher dimensions, it is successful for some applications in quantum field theory. In particular, here we restrict ourselves to the Gaussian setting, which corresponds to the Gaussian Free Field model for which MGMC has been shown to be an efficient sampler in [undef]. On the other hand, the way our Alg. 1 is written down makes it explicit that the linear operators for transferring samples between levels of the hierarchy are not unique. This allows a generalisation of the MGMC algorithm to algebraic multigrid (AMG) settings or more sophisticated FE discretisations. While arbitrary intergrid operators are also implicitly included in the algorithms presented in [undef], the authors of that paper focus on the case of piecewise linear interpolation for finite difference discretisations on regular meshes. As in the standard multigrid algorithm for solving linear systems, we introduce a prolongation matrix $I_{\ell-1}^{\ell}$ defined in (2.14); this matrix has full column rank for all levels $\ell$. The corresponding restriction matrix $I_{\ell}^{\ell-1}=(I_{\ell-1}^{\ell})^{\top}$ is the transpose of the prolongation matrix. Coarse level matrices are then recursively constructed via the so-called Galerkin triple-product
@@ 3.1.1 Algorithm description
9:   Define $f_{\ell-1}:=I_{\ell}^{\ell-1}(f_{\ell}-\widetilde{A}_{\ell}\theta_{\ell,\nu_{1}})$ $\triangleright$ Restriction
@@
12:     $\psi_{\ell-1}^{(m+1)}:=\mathrm{MGMC}_{\ell-1}(\widetilde{A}_{\ell-1},f_{\ell-1},\psi_{\ell-1}^{(m)})$ $\triangleright$ Recursive call to $\mathrm{MGMC}_{\ell-1}$
@@
14:   $\theta_{\ell,\nu_{1}+1}:=\theta_{\ell,\nu_{1}}+I_{\ell-1}^{\ell}\psi_{\ell-1}^{(\gamma_{\ell})}$ $\triangleright$ Prolongation
@@
16:     $\theta_{\ell,\nu_{1}+2+j}:=\mathrm{RandomSmoother}(\widetilde{A}_{\ell},\widetilde{M}_{\ell}^{\mathrm{post}},f_{\ell},\theta_{\ell,\nu_{1}+1+j})$ $\triangleright$ Random post-smoothing (Alg. 2)
@@
    If the initial state $\theta_{L}^{0}\sim\mathcal{N}(\mu_{L}^{0},(\widetilde{A}_{L}^{0})^{-1})$ is drawn from a multivariate normal distribution with mean $\mu_{L}^{0}=(\widetilde{A}_{L}^{0})^{-1}f_{L}^{0}$ and covariance $(\widetilde{A}_{L}^{0})^{-1}$, the Markov chain $\theta_{L}^{(0)},\theta_{L}^{(1)},\theta_{L}^{(2)},\dots$ generated by $\textrm{MGMC}_{L}(\widetilde{A}_{L},f_{L},\cdot)$ converges to the target distribution $\mathcal{N}(\mu_{L},\widetilde{A}_{L}^{-1})$ in the sense that $\theta_{L}^{(k)}\sim\mathcal{N}(\mu_{L}^{k},(\widetilde{A}_{L}^{k})^{-1})$ with $\mu_{L}^{k}\rightarrow\mu_{L}$ and $\widetilde{A}_{L}^{k}\rightarrow\widetilde{A}_{L}$. The rate of convergence is grid-independent for both the mean and the covariance (see Section 4.3).
@@
The intuitive explanation for the success of MGMC is the same as for multigrid solvers: by using a hierarchy of levels, the samples are updated on all length scales simultaneously and this incurs only a small, grid independent overhead. This is in stark contrast to a Gibbs sampler, which only updates the samples locally. The same idea can also be understood spectrally: a sample $\theta_{L}^{(k)}$ in the Markov chain can be expanded as
$\theta_{L}^{(k)}=\mu_{L}+\sum_{j=1}^{n_{L}}\widetilde{v}_{L,j}\widetilde{\lambda}_{L,j}^{-1/2}\xi_{L,j}^{(k)}$ (3.2)
where $\widetilde{\lambda}_{L,j}$ and $\widetilde{v}_{L,j}$ are the eigenvalues and normalised eigenvectors of $\widetilde{A}_{L}$ respectively and for samples that are drawn from the target distribution $\xi_{L,j}^{(k)}\sim\mathcal{N}(0,1)$ are i.i.d. Gaussian random variables. Small eigenvalues correspond to smooth eigenvectors that vary slowly in space. While one step of the Gibbs-sampler only significantly changes the spectral expansion coefficients $\xi_{L,j}^{(k)}$ that correspond to high-frequency eigenvectors, MGMG will modify all expansion coefficients simultaneously for suitable operators $\mathcal{A}$. This should be compared to the action of a deterministic multigrid update which reduces the error on all length scales, i.e. for all eigenvectors.